# TRENDY: Gene Regulatory Network Inference Enhanced by Transformer

### v1

@@ Abstract
Gene regulatory networks (GRNs) play a crucial role in the control of cellular functions. Numerous methods have been developed to infer GRNs from gene expression data, including mechanism-based approaches, information-based approaches, and more recent deep learning techniques, the last of which often overlooks the underlying gene expression mechanisms. In this work, we introduce TRENDY, a novel GRN inference method that integrates transformer models to enhance the mechanism-based WENDY approach. Through testing on both simulated and experimental datasets, TRENDY demonstrates superior performance compared to existing methods. Furthermore, we apply this transformer-based approach to three additional inference methods, showcasing its broad potential to enhance GRN inference.
@@ 1 Introduction
The expression of genes can be regulated by other genes. We can construct a directed graph, where vertices are genes, and edges are regulatory relationships. This graph is called a gene regulatory network (GRN). GRNs are essential to understanding the mechanisms governing cellular processes, including how cells respond to various stimuli, differentiate, and maintain homeostasis [1, 2]. Understanding the GRN structure is important for developmental biology [3, 4, 5, 6], disease research [7, 8, 9], and even studying macroscopic behavior [10, 11, 12].
It is difficult to directly determine the GRN structure with experiments. Instead, researchers develop methods that infer the GRN structure through gene expression data. Some methods, such as WENDY [13] and NonlinearODEs [14], are mechanism-based: they construct dynamical models for gene expression, and fit with expression data to determine the GRN [15, 16]. Some methods, like GENIE3 [17] and SINCERITIES [18], are information-based: they directly find genes that contain information (predictability) for the target gene, and assume that such information implies regulation [19, 20].
Recently, there are some deep learning-based methods that infer GRN with different neural network frameworks, such as convolutional neural networks [21], recurrent neural networks [22], variational autoencoders [23], graph neural networks [24, 25], etc. Deep learning frameworks can perform well on many tasks. However, because of the large number of tunable parameters in a neural network, it is difficult to explain why it works. These deep learning-baed GRN inference methods have similar problems: they generally apply popular neural networks as black boxes without integration with biological mechanisms, and thus lacking interpretability.
The transformer model is a deep learning architecture designed to handle sequential data [26]. It can capture long-range dependencies and relationships through the self-attention mechanism. The encoder layer of the transformer model can process the input through a neural network, so that the output is close to the given target.
@@
To combine the powerful deep learning techniques and biological understanding of gene regulation, we propose a novel GRN inference method that applies transformer models to enhance a mechanism-based GRN inference method, WENDY. This new method, TRansformer-Enhanced weNDY, is abbreviated as TRENDY. WENDY method is based on a dynamical model for the covariance matrix of genes. In TRENDY, we first use a transformer model to construct a pseudo-covariance matrix that performs better in WENDY. Then we apply another transformer model that directly enhances the inferred GRN.
The idea for the second half of TRENDY can be used to enhance the inferred results by other GRN inference methods. We apply transformer models to three GRN inference methods, GENIE3, SINCERITIES, and NonlinearODEs, to obtain their enhanced versions, GENIE3-rev, SINCERITIES-rev, and NonlinearODEs-rev.
We test four transformer-enhanced methods against their non-transformer counterparts on two simulated data sets and two experimental data sets. All four transformer-enhanced methods outperform existing methods, and TRENDY ranks the first among all methods.
@@
The matrix $A_{\mathrm{true}}$ is a randomly generated ground truth GRN, where $(A_{\mathrm{true}})_{i,j}>0/=0/<0$ means that gene $i$ has positive/no/negative regulatory effects on gene $j$. $X_{i}(t)$ is the expression level of gene $i$ at time $t$, and $W_{j}(t)$ is a standard Brownian motion, independent with other $W_{i}(\tau)$. The parameter values are $V=30$, $\beta=1$, $\theta=0.2$, and $\sigma=0.1$. Our goal is to infer $A_{\mathrm{true}}$ from $X_{j}(t)$.
@@ 2.1 TRENDY method
meaning that $(I+tA_{\mathrm{true}}^{\mathrm{T}})K_{0}(I+tA_{\mathrm{true}})$ and $(I+tA_{0}^{\mathrm{T}})K_{0}(I+tA_{0})$ differ by $E$, the error introduced by the linearization, and $A_{0}$ may not match $A_{\mathrm{true}}$ accurately. If we want to derive a more accurate and complicated equation of $K_{0}$, $K_{t}$, and $A_{\mathrm{true}}$ than Eq. 2, this equation might be numerically unstable to solve. Therefore, we need to find another way to improve the WENDY method.
@@
Fig. 1 describes the training and testing workflow for TRENDY. To train TRENDY, one needs to
(1) generate many random $A_{\mathrm{true}}$ and their corresponding gene expression data;
(2) calculate the covariance matrices $K_{0}$ and $K_{t}$, and then calculate $K_{t}^{*}$ from $K_{0},K_{t},A_{\mathrm{true}}$;
(3) use $K_{t}$ as input and $K_{t}^{*}$ as target to train the $\mathrm{TE}(k=1)$ model, and then use the trained $\mathrm{TE}(k=1)$ model to calculate $K_{t}^{\prime}$ from $K_{t}$;
(4) use WENDY to calculate $A_{1}$ from $K_{0},K_{t}^{\prime}$;
(5) use $A_{1},K_{0},K_{t}$ as input and $A_{\mathrm{true}}$ as target to train the $\mathrm{TE}(k=3)$ model.
For applying TRENDY to infer GRN from existing data, one needs to omit step (1) and
(2) calculate the covariance matrices $K_{0}$ and $K_{t}$;
(3) use the trained $\mathrm{TE}(k=1)$ model to calculate $K_{t}^{\prime}$ from $K_{t}$;
(4) use WENDY to calculate $A_{1}$ from $K_{0},K_{t}^{\prime}$;
(5) feed $A_{1},K_{0},K_{t}$ to the trained $\mathrm{TE}(k=3)$ model to calculate the final inferred GRN $A_{2}$, the final output of TRENDY.
See https://github.com/YueWangMathbio/TRENDY for all code files and data used in this paper.
@@ 2.2 Transformer model $\mathrm{TE}(k)$
We build a deep learning framework $\mathrm{TE}(k)$ that is based on transformer encoder layers, where $k$ is the major hyperparameter that describes the number of input matrices. Besides, there are three hyperparameters that can be tuned: the model dimension $d$; the number of encoder layers $l$; the number of attention heads $h$ of the encoder layer. We use this $\mathrm{TE}(k)$ model with different hyperparameters in TRENDY and other transformer-enhanced methods. See Table 1 for the general structure of $\mathrm{TE}(k)$, along with the shape of data after each layer. Notice that the same transformer encoder layer can handle inputs of different lengths. This means that the gene number $n$ is not a predetermined hyperparameter, and we do not need to train a different model for each $n$. We will train the $\mathrm{TE}(k)$ model with $n=10$ genes and test on data sets with $n=10/18/20$ genes.
@@
input ($k$ matrices) $k$ groups of $n\times n$
linear embedding layer ($1$ to $d$) $k$ groups of $n\times n\times d$
segment embedding layer (omitted when $k=1$) $k$ groups of $n\times n\times d$
2-D positional encoding layer $k$ groups of $n\times n\times d$
flattening and concatenation $(n^{2})\times(dk)$
$l$ layers of transformer encoder with head number $h$ $(n^{2})\times(dk)$
linear embedding layer ($(dk)$ to $1$) $n^{2}\times 1$
reshaping and output $n\times n$
training target $n\times n$
Table 1: Structure of the $\mathrm{TE}(k)$ model. The left column lists all layers. The right column illustrates the shape of data after each layer, where $n$ is the number of genes, $d$ is the model dimension, and $k$ is the number of input matrices. Notice that $k,d,l,h$ are tunable hyperparameters, but $n$ is not a predetermined hyperparameter.
@@ 2.3 Other transformer-enhanced methods
SINCERITIES [18] works on single-cell expression data at multiple time points. It runs regression on the difference between distributions of expression levels, also being information-based.
@@
For inferred GRN $A_{\mathrm{G}}$ by GENIE3, we train a $\mathrm{TE}(k=2)$ model with $A_{\mathrm{G}}$ and the covariance matrix $K$ as input. Then we can use this trained model to calculate a more accurate $A_{\mathrm{G}}^{\prime}$ from $A_{\mathrm{G}}$ and $K$. This method is named GENIE3-rev.
For inferred GRN $A_{\mathrm{S}}$ by SINCERITIES, we train a $\mathrm{TE}(k=1)$ model with $A_{\mathrm{S}}$ as input, and name it SINCERITIES-rev.
For inferred GRN $A_{\mathrm{N}}$ by NonlinearODEs, we train a $\mathrm{TE}(k=1)$ model with $A_{\mathrm{N}}$ as input, and name it NonlinearODEs-rev.
@@ 2.4 Performance of different methods
The same as the WENDY paper [13], we test the performance of different GRN inference methods on two simulated data sets and two experimental data sets. To compare the inferred GRN $A_{\mathrm{pred}}$ and the ground truth GRN $A_{\mathrm{true}}$, we adopt the common practice that calculates two area under curve (AUC) scores: AUROC and AUPRC [13]. Both AUC scores are between 00 and $1$, where $1$ means perfect matching, and 00 means perfect mismatching.
2.4.1 SINC data sets
We use the same method for generating training data to generate 1000 testing samples, and name this data set as SINC [18, 13]. Each sample with a random $A_{\mathrm{true}}$ has 100 cells measured at time points $0.0,0.1,\ldots,1.0$.
For WENDY and TRENDY, we set the first time point to $0.0$ and vary the second time point from $0.1$ to $1.0$. For GENIE3 and GENIE3-rev, we set the time point to any of $0.1,\ldots,1.0$. For SINCERITIES, SINCERITIES-rev, NonlinearODEs, and NonlinearODEs-rev, we use all time points. See Fig. 2 and Table 2 for the results.
Figure 2: Average AUROC and AUPRC values of eight GRN inference methods on the SINC data set. Notice that SINCERITIES, NonlinearODEs and their revised versions need data at all time points.
2.4.2 DREAM4 data set
DREAM4 data set is also generated by simulating an SDE system, and consists of multiple data sets under different settings [39]. DREAM4 is commonly used as a benchmark for GRN inference [13, 18]. Here we use the 10-gene time series data set. There are five $A_{\mathrm{true}}$, each corresponding to the levels of 10 genes for five cells at 21 time points.
For WENDY and TRENDY, we consider any pair of time points and take average. For GENIE3 and GENIE3-rev, we consider any one time point and take average. For SINCERITIES, SINCERITIES-rev, NonlinearODEs, and NonlinearODEs-rev, we use any consecutive 11 time points and take average. See Fig. 3 and Table 2 for the results.
Figure 3: Average AUROC and AUPRC values of eight GRN inference methods on the DREAM4 data set.
2.4.3 THP-1 data set
THP-1 data set measures monocytic THP-1 human myeloid leukemia cells at eight time points [40]. Each time point has 120 cells measured. The same as previous papers [18, 13], we consider the expression levels of 20 genes.
For WENDY and TRENDY, we consider any pair of time points and take average. For GENIE3 and GENIE3-rev, we consider any one time point and take average. For SINCERITIES, SINCERITIES-rev, NonlinearODEs, and NonlinearODEs-rev, we use all eight time points. See Fig. 4 and Table 2 for the results.
Figure 4: AUROC and AUPRC values of eight GRN inference methods on the THP-1 data set.
2.4.4 hESC data set
hESC data set measures human embryonic stem cell-derived progenitor cells at six time points [41]. Each time point has 66–172 cells measured. The same as in a previous paper [13], we consider the expression levels of 18 genes.
For WENDY and TRENDY, we consider any pair of time points and take average. For GENIE3 and GENIE3-rev, we consider any one time point and take average. For SINCERITIES, SINCERITIES-rev, NonlinearODEs, and NonlinearODEs-rev, we use all six time points. See Fig. 5 and Table 2 for the results. Notice that the ground truth GRN $A_{\mathrm{true}}$ has very few nonzero values. This is why the AUPRC value is much lower than the AUROC value.
Figure 5: AUROC and AUPRC values of eight GRN inference methods on the hESC data set.
2.4.5 Overall evaluation
@@
We summarize the performance of all eight methods on four data sets. See Table 2 for details. The four transformer-enhanced methods (TRENDY, GENIE3-rev, SINCERITIES-rev, NonlinearODEs-rev) are all better than the four known methods. This means that our idea of enhancing GRN inference methods with $\mathrm{TE}(k)$ models works well universally. Although the training data only has $n=10$ genes, and the THP-1 and hESC data sets have $n=20$ and $n=18$ genes, the overall performance of transformer-enhanced methods is better than their non-transformer counterparts.
@@
The core method of this paper, TRENDY, ranks the first among all eight methods. Besides, the TRENDY method is better than the WENDY method on three data set, and has almost the same performance as WENDY on one data set (THP-1). The other three transformer-enhanced methods are worse than their non-transformer counterparts on at least one data set. This means that TRENDY’s performance is robust on different data sets. In sum, TRENDY has satisfactory performance among all methods tested.
WEN TRE GEN GENr SIN SINr NON NONr
S-RO 0.6654 0.8547 0.4062 0.8703 0.6783 0.7964 0.5076 0.5976
S-PR 0.6109 0.7519 0.5061 0.7672 0.5829 0.6637 0.5313 0.5658
D-RO 0.4899 0.5341 0.5636 0.4589 0.4908 0.4995 0.4806 0.5712
D-PR 0.2080 0.2177 0.2286 0.1799 0.1919 0.2034 0.1705 0.2452
T-RO 0.5261 0.5557 0.4484 0.5506 0.6261 0.5251 0.5338 0.4808
T-PR 0.3972 0.3669 0.3546 0.3781 0.3852 0.3412 0.3486 0.3302
h-RO 0.4997 0.5311 0.5913 0.6008 0.4198 0.4871 0.5971 0.6233
h-PR 0.0392 0.0376 0.0468 0.0435 0.0261 0.0294 0.0534 0.0641
Total 3.4364 3.8497 3.1456 3.8493 3.4011 3.5458 3.2229 3.4782
Rank 5th 1st 8th 2nd 6th 3rd 7th 4th
Table 2: Performance summary of eight GRN inference methods on four data sets. For row names, S, D, T, h stand for SINC, DREAM4, THP-1, hESC data sets; -RO and -PR represent AUROC and AUPRC values. For column names, WEN, TRE, GEN, GENr, SIN, SINr, NON, NONr represent WENDY, TRENDY, GENIE3, GENIE3-rev, SINCERITIES, SINCERITIES-rev, NonlinearODEs, NonlinearODEs-rev. The “Total” row is the sum of all eight AUC values, and the last row is the rank of each method by the total score. Red numbers indicate that they are smaller than their non-transformer counterparts (left neighbors) by at least 0.01, and blue color means larger by at least 0.01. For instance, TRENDY has S-RO$=0.8547$, larger than WENDY’s S-RO$=0.6654$ by at least 0.01, so that $0.8547$ is in blue.

### v2

@@ Abstract
Gene regulatory networks (GRNs) play a crucial role in the control of cellular functions. Numerous methods have been developed to infer GRNs from gene expression data, including mechanism-based approaches, information-based approaches, and more recent deep learning techniques, the last of which often overlook the underlying gene expression mechanisms. In this work, we introduce TRENDY, a novel GRN inference method that integrates transformer models to enhance the mechanism-based WENDY approach. Through testing on both simulated and experimental datasets, TRENDY demonstrates superior performance compared to existing methods. Furthermore, we apply this transformer-based approach to three additional inference methods, showcasing its broad potential to enhance GRN inference.
@@ 1 Introduction
The expression of genes can be regulated by other genes. We can construct a directed graph, where vertices are genes, and edges are regulatory relationships. This graph is called a gene regulatory network (GRN). GRNs are essential to understanding the mechanisms governing cellular processes, including how cells respond to various stimuli, differentiate, and maintain homeostasis [1, 2]. Understanding the GRN structure is important for developmental biology [3, 4, 5, 6, 7], disease research [8, 9, 10, 11], and even studying macroscopic behavior [12, 13, 14].
It is difficult to directly determine the GRN structure with experiments. Instead, researchers develop methods that infer the GRN structure through gene expression data. Some methods, such as WENDY [15] and NonlinearODEs [16], are mechanism-based: they construct dynamical models for gene expression, and fit with expression data to determine the GRN [17, 18]. Some methods, like GENIE3 [19] and SINCERITIES [20], are information-based: they treat this as a feature selection problem and directly find genes that can be used to predict the level of the target gene (e.g., by linear regression). Then this predictability (information) for the target gene is assumed to imply regulation [21, 22].
Recently, there are some deep learning-based methods that infer GRN with different neural network frameworks, such as convolutional neural networks [23], recurrent neural networks [24], variational autoencoders [25], graph neural networks [26, 27], etc. Deep learning frameworks can perform well on many tasks. However, because of the large number of tunable parameters in a neural network, it is difficult to explain why it works. These deep learning-based GRN inference methods have similar problems: they generally apply popular neural networks as black boxes without integration with biological mechanisms, and thus lacking interpretability.
The transformer model is a deep learning architecture designed to handle sequential data [28]. It can capture long-range dependencies and relationships (linear and nonlinear) through the self-attention mechanism. The encoder layer of the transformer model can process the input through a neural network, so that the output is close to the given target.
@@
To combine the powerful deep learning techniques and biological understanding of gene regulation, we propose a novel GRN inference method that applies transformer models to enhance a mechanism-based GRN inference method, WENDY. This new method, TRansformer-Enhanced weNDY, is abbreviated as TRENDY. WENDY method is based on a dynamical model for gene regulation, and an equation for the GRN and the covariance matrix of genes is solved to derive the GRN. In TRENDY, we first use a transformer model to construct a pseudo-covariance matrix that performs better in WENDY. Then we apply another transformer model that directly enhances the inferred GRN.
The idea for the second half of TRENDY can be used to enhance the inferred results by other GRN inference methods. We apply transformer models to three GRN inference methods, GENIE3, SINCERITIES, and NonlinearODEs, to obtain their transformer-enhanced versions, tGENIE3, tSINCERITIES, and tNonlinearODEs.
There have been some research papers that study the idea of denoising networks (e.g., enhancing GRNs inferred by other methods), such as network deconvolution (ND) [32], diffusion state distance [33], BRANE Cut (BC) [34], BRANE Clust [35], and network enhancement [36].
We test the four traditional methods (WENDY, GENIE3, SINCERITIES, and NonlinearODEs) with their enhanced versions (by transformer, ND, or BC) on two simulated data sets and two experimental data sets. All four transformer-enhanced methods outperform other methods, and TRENDY ranks the first among all methods.
@@
The matrix $A_{\mathrm{true}}$ is a randomly generated ground truth GRN, where $(A_{\mathrm{true}})_{i,j}>0/=0/<0$ means that gene $i$ has positive/no/negative regulatory effects on gene $j$. $X_{i}(t)$ is the expression level of gene $i$ at time $t$, and $W_{j}(t)$ is a standard Brownian motion. For different genes, the corresponding Brownian motions are independent. The same as in previous papers [43, 20, 15], the parameter values are $V=30$, $\beta=1$, $\theta=0.2$, and $\sigma=0.1$. Our goal is to infer $A_{\mathrm{true}}$ from $X_{j}(t)$.
@@ 2.1 TRENDY method
meaning that $(I+tA_{\mathrm{true}}^{\mathrm{T}})K_{0}(I+tA_{\mathrm{true}})$ and $(I+tA_{0}^{\mathrm{T}})K_{0}(I+tA_{0})$ differ by $E$, the error introduced by the linearization, and $A_{0}$ may not match $A_{\mathrm{true}}$ accurately. If we want to derive a more accurate and complicated equation of $K_{0}$, $K_{t}$, and $A_{\mathrm{true}}$ than Eq. 2, this equation might be numerically unstable to solve. Therefore, we need to find another way to improve WENDY.
@@
The actual $K_{t}$ is a complicated nonlinear function of $A_{\mathrm{true}}$ and $K_{0}$. Thus $K_{t}^{*}$, also a function of $A_{\mathrm{true}}$ and $K_{0}$, can be roughly learned by the $\mathrm{TE}(k=1)$ model as a function of $K_{t}$. Given accurate $K_{0}$ and $K_{t}^{*}$, we still cannot uniquely determine $A_{\mathrm{true}}$, and the numerical solver for Eq. 3 determines which $A_{1}$ is chosen. Therefore, the $\mathrm{TE}(k=3)$ model can partially learn the nonlinear behavior of the solver and output a better $A_{2}$ from $A_{1}$.
Fig. 1, Algorithm 1, and Algorithm 2 describe the training and testing workflow for TRENDY.
  1. 1.
    Repeat generating random $A_{\mathrm{true}}$ and corresponding gene expression data
  2. 2.
    Calculate covariance matrices $K_{0}$ and $K_{t}$, and then calculate $K_{t}^{*}$ from $K_{0},K_{t},A_{\mathrm{true}}$
  3. 3.
    Train $\mathrm{TE}(k=1)$ model with input $K_{t}$ and target $K_{t}^{*}$
    Call trained $\mathrm{TE}(k=1)$ model to calculate $K_{t}^{\prime}$ from $K_{t}$
  4. 4.
    Call WENDY to calculate $A_{1}$ from $K_{0},K_{t}^{\prime}$
  5. 5.
    Train $\mathrm{TE}(k=3)$ model with input $A_{1},K_{0},K_{t}$ and target $A_{\mathrm{true}}$
Algorithm 1 Training workflow of TRENDY method.
  1. 1.
    Input: gene expression data at two time points
  2. 2.
    Calculate covariance matrices $K_{0}$ and $K_{t}$
  3. 3.
    Call trained $\mathrm{TE}(k=1)$ model to calculate $K_{t}^{\prime}$ from $K_{t}$
  4. 4.
    Call WENDY to calculate $A_{1}$ from $K_{0},K_{t}^{\prime}$
  5. 5.
    Call trained $\mathrm{TE}(k=3)$ model to calculate $A_{2}$ from $A_{1},K_{0},K_{t}$
  6. 6.
    Output: inferred GRN $A_{2}$
Algorithm 2 Testing workflow of TRENDY method.
@@ 2.2 Transformer model $\mathrm{TE}(k)$
We build a deep learning framework $\mathrm{TE}(k)$ that is based on transformer encoder layers, where $k$ is the major hyperparameter that describes the number of input matrices. Besides, there are three hyperparameters that can be tuned: the model dimension $d$; the number of encoder layers $l$; the number of attention heads $h$ of the encoder layer. We use this $\mathrm{TE}(k)$ model with different hyperparameters in TRENDY and other transformer-enhanced methods. See Algorithm 3 for the general structure of $\mathrm{TE}(k)$, along with the shape of data after each layer. Besides the standard transformer structure, the segment embedding layer tells the model to treat $k>1$ input matrices separately, and the 2-D positional encoding layer tells the model to remember that the input is originally two-dimensional. Notice that the same transformer encoder layer can handle inputs of different lengths. This means that the gene number $n$ is not a predetermined hyperparameter, and we do not need to train a different model for each $n$. We will train the $\mathrm{TE}(k)$ model with $n=10$ genes and test on data sets with $n=10/18/20$ genes.
@@
  1. 1.
    Input: $k$ matrices ($k$ groups of $n\times n$)
  2. 2.
    Linear embedding layer with dimension $d$ ($k$ groups of $n\times n\times d$)
  3. 3.
    Segment embedding layer (omitted when $k=1$) ($k$ groups of $n\times n\times d$)
  4. 4.
    2-D positional encoding layer ($k$ groups of $n\times n\times d$)
  5. 5.
    Flattening and concatenation ($(n^{2})\times(dk)$)
  6. 6.
    $l$ layers of transformer encoder ($(n^{2})\times(dk)$)
  7. 7.
    Linear embedding layer ($n^{2}\times 1$)
  8. 8.
    Output: reshaping into a matrix ($n\times n$)
Algorithm 3 Structure of the $\mathrm{TE}(k)$ model. The shape of data after each layer is in the brackets.
@@ 2.3 Other transformer-enhanced methods
The second half of the TRENDY method applies a transformer model to learn the highly nonlinear relation between $A_{1}$ (result of WENDY) and $A_{\mathrm{true}}$. For other GRN inference methods, we can assume that the inferred GRN and $A_{\mathrm{true}}$ have some nonlinear and possibly random relations, since these methods only consider and infer certain forms of regulations. Then we can try to enhance the inferred results by a transformer model.
@@
SINCERITIES [20] works on single-cell expression data at multiple time points. It runs regression on the Kolmogorov–Smirnov distance between cumulative distribution functions of expression levels, also being information-based.
@@
For inferred GRN $A_{\mathrm{G}}$ by GENIE3, we train a $\mathrm{TE}(k=2)$ model with $A_{\mathrm{G}}$ and the covariance matrix $K$ as input. Then we can use this trained model to calculate a more accurate $A_{\mathrm{G}}^{\prime}$ from $A_{\mathrm{G}}$ and $K$. This method is named tGENIE3.
For inferred GRN $A_{\mathrm{S}}$ by SINCERITIES, we train a $\mathrm{TE}(k=1)$ model with $A_{\mathrm{S}}$ as input, and name it tSINCERITIES.
For inferred GRN $A_{\mathrm{N}}$ by NonlinearODEs, we train a $\mathrm{TE}(k=1)$ model with $A_{\mathrm{N}}$ as input, and name it tNonlinearODEs.
2.4 Methods enhanced by ND and BC
Network deconvolution (ND) [32] requires the input to be nonnegative and symmetric, so as its output. The idea is to assume that the input is a convolution of a simpler network, and solves this simpler network by deconvolution. BRANE Cut (BC) [34] performs well when we know that some genes are transcription factors and others are not. When we have no knowledge of transcription factors, BC degenerates to setting a threshold for whether one regulation exists. BC requires the input to be nonnegative and symmetric, and the output is a 0-1 matrix. We use ND and BC to enhance those four traditional methods, and name them as nWENDY, nGENIE3, nSINCERITIES, nNonlinearODEs, bWENDY, bGENIE3, bSINCERITIES, bNonlinearODEs.
@@ 2.5 Performance of different methods
Certain autoregulations can decrease the variance [46, 47, 48], and mRNA bursts can increase the gene expression variance [49, 50, 51], which can be approximated by tuning the value of $\sigma$ in Eq. 1. SINC data set is generated by Eq. 1 with $\sigma=0.01/0.1/1$. The synthetic DREAM4 data set [52] is commonly used as a benchmark for GRN inference [15, 20]. THP-1 data set measures monocytic THP-1 human myeloid leukemia cells [53]. hESC data set measures human embryonic stem cell-derived progenitor cells[54].
@@
For these four data sets, two synthetic and two experimental, we test 16 methods: four traditional methods (WENDY, GENIE3, SINCERITIES, and NonlinearODEs), their transformer variants, ND variants, and BC variants. To compare the inferred GRN $A_{\mathrm{pred}}$ and the ground truth GRN $A_{\mathrm{true}}$, we adopt the common practice that calculates two area under curve (AUC) scores: AUROC and AUPRC [15]. Both AUC scores are between 00 and $1$, where $1$ means perfect matching, and 00 means perfect mismatching. See Section S2 for details of these two scores.
Figure 2: AUROC and AUPRC scores of WENDY and its variants on four data sets.
Figure 3: AUROC and AUPRC scores of GENIE3 and its variants on four data sets.
Figure 4: AUROC and AUPRC scores of SINCERITIES and its variants on four data sets.
Figure 5: AUROC and AUPRC scores of NonlinearODEs and its variants on four data sets.
Figure 6: Total scores of all 16 methods, where the $x$-axis starts at 2.8.
@@
See Figs. 26 and Tables S1,S2 for the performance of these methods. Although the training data only has $n=10$ genes, and the THP-1 and hESC data sets have $n=20$ and $n=18$ genes, the four transformer-enhanced methods (TRENDY, tGENIE3, tSINCERITIES, tNonlinearODEs) are better than all other methods. This means that our idea of enhancing GRN inference methods with $\mathrm{TE}(k)$ models works well universally. ND and BC generally do not enhance the original method significantly.
@@
The core method of this paper, TRENDY, ranks the first among all 16 methods (tGENIE3 is slightly inferior). Besides, the TRENDY method is better than the WENDY method on three data set, and has almost the same performance as WENDY on one data set (THP-1). The other three transformer-enhanced methods are worse than their non-transformer counterparts on at least one data set. This means that TRENDY’s performance is robust on different data sets. In sum, TRENDY has satisfactory performance among all methods tested.
All methods have very small AUPRC values on hESC data set. The reason is that the ground truth GRN $A_{\mathrm{true}}$ has very few nonzero values (Fig. S3).